A ray of light is plane polarized. A sheer of Polaroid is held at an angle of 60 to the rat. What fraction of the original intensity if the ray is transmitted through the Polaroid at an angle of 60.
1
Expert's answer
2020-12-10T11:06:41-0500

Fraction of the original intensity "\eta =I/I_0" transmitted through the Polaroid is given by the Malus's law (see https://en.wikipedia.org/wiki/Polarizer#Malus.27s_law_and_other_properties):


"\eta = \dfrac{I}{I_0} = \cos^2\alpha"

where "I_0" is the initial intencity, "I" is the transmitted intencity, and "\alpha =60\degree" is the angle of the polaroid. Obtain:


"\eta = \cos^260\degree = \dfrac{1}{4} = 0.25"

Answer. 0.25.